OpenStudy (anonymous):

Could you please help me with this problem: 5w+3z-10+w solve for z

OpenStudy (anonymous):

Do you have a given w value or another equation? If not, then z must be solved for simply in terms of w. So, you get: 5w+3z-10+w = 6w+3z-10 3z = 10-6w z = (10/3)-2w

OpenStudy (anonymous):

I have a list of answers because it was a multiple choice test earlier today. I handed in my scantron but was allowed to take the test section home. I am trying to learn the ones I had problems with. That answer is not on my sheet. Is there another way to solve it?